FTIR MATRIX STUDY OF POTENTIAL CIRCUMSTELLAR MOLECULES: TiC_{3}

\underline{R. E. KINZER, JR.}, C. M. L. RITTBY, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

Results will be presented of recently initiated studies on the structures and infrared spectra of transition metal-carbon clusters that may be of interest in circumstellar shells or other astrophysical environments. The FTIR (Fourier transform infrared) spectrum of TiC_{3} was observed by trapping the vapor produced during dual Nd:YAG laser ablation of Ti and C rods in solid Ar at ∼10 K. Measurements of frequencies and ^{13}C isotopic shifts have enabled the identification of the fan-like (C_{2v}) isomer of TiC_{3} with vibrational fundamentals ν_{3}(a_{1}) = 624.3 and ν_{5}(b_{2}) = 1484.2 cm^{−1}. The results are in good agreement with the predictions of DFT calculations at the B3LYP/6-311G(3df, 3pd) level. The observed C_{2v} structure is also consistent with the results from an earlier photoelectron spectroscopy study.^{a}

aX. B. Wang, C. F. Ding, L. S. Wang, J. Phys. Chem. A \underline{101}, 7699 (1997).